# LUCI in the surface code with dropouts

### v1

@@ III Results
In this section we will compare LUCI to the two other methods of handling dropout mentioned in the introduction. We will refer to the method from Auger et al. (2017) and Wei et al. (2024) as the Auger/Bandage method, and the method from Strikis et al. (2023); Siegel et al. (2023); Lin et al. (2024) as the Strikis/Brown method. As shown in Wei et al. (2024), the Strikis/Brown method removes additional qubits and is strictly worse than the Auger/Bandage method, but is helpful because the single-type boundaries make certain proofs more straightforward. We keep it as a comparison point since it is well known, and as a result serves as a point of reference. We use stim for all simulations, with a correlated minimum-weight perfect matching decoder used for decoding and SI1000 noise Gidney et al. (2021), which sets error rates for different operations based on a single parameter $p$, described in Appendix A. We simulate X and Z logical memory experiments for ××$\ell\times\ell\times 4\ell$ code blocks, where $\ell$ is the patch diameter of the grid, and convert the resulting average logical error to a per-round logical error rate. For LUCI circuits we define a round as a single round of the LUCI diagram, meaning that we will use each round $\ell$ times over the course of the $4\ell$ rounds. As these different methods have different timelike distance, this comparison is slightly complicated, as depending on the timelike distance of your circuit you may need more rounds. LUCI circuits only promise to measure all stabilizers over four rounds, while the other two methods measure all stabilizers in two. We note, however, that the LUCI circuits still measure the majority of stabilizers at the same rate as standard methods, and only at low error rates will the sparse low-weight logical operators dominate performance. In practice, one should run stability experiments to chose the number of rounds in a logical idle block Gidney (2022). The circuits used are provided in Debroy (2024).
@@
For the Auger/Bandage method and LUCI, we use custom written circuit generation tools, while for the Strikis/Brown method we use the circuit generation software helpfully provided in Lin et al. (2024). These methods are both amenable to using shells, first described in Higgott and Breuckmann (2021) as gauge-fixings, and then used in Strikis et al. (2023) to prove the existence of thresholds for surface codes under dropouts. In our comparison we do not use shells, but Wei et al. (2024) discusses how they improve performance for the bandage method, and we expect the same would hold for LUCI. Appendix E discusses how one could implement shells with LUCI.
@@
In Fig. 9, we compare the three methods on the example grid presented in Fig. 1. The LUCI circuit used is the one built in Figs. 7 and 8. By looking at the slopes of the three curves, along with the gray reference curve, we can see the impact of distance loss as the error rate improves. The baseline distance-9 circuit has a slope of 5 in the log-log plot, because it takes $\lfloor\frac{d+1}{2}\rfloor$ errors to cause a logical error. When dropouts are added into the device, the distance is reduced to 3 for the Strikis/Brown and Auger/Bandage methods, and 7 for LUCI. The difference between the Strikis/Brown and Auger/Bandage methods is around the broken data qubit and broken coupler labeled as (e) in Fig. 1. As explained in Ref. Wei et al. (2024), this configuration produces a “bridge” qubit, which the Auger/Bandage method preserves and the Strikis/Brown method removes. Otherwise the two circuits are very similar, and perform near identically as seen in the figure. LUCI also keeps this qubit, and only removes an additional qubit for the two-broken-coupler and broken-coupler-near-broken-qubit configurations labeled by (d) and (f) in Fig. 1. At an error rate of 0.001, LUCI shows more than a 30×30$\times$30 × improvement in logical error rate per round over Auger/Bandage and Strikis/Brown methods for this example grid.
@@
Figure 11: Teraquop footprint plot for LUCI. 100 grids were sampled for each error rate and dropout rate at $\ell\in{7,11,15}$ and then projected to $10^{-12}$. Error rates are for codeblocks of size $\ell\times\ell\times\ell$ for the dropout-free baseline circuits, ××$\ell\times\ell\times 2\ell$ for the Strikis/Brown and Auger/Bandage circuits, and ××$\ell\times\ell\times 4\ell$ for the LUCI circuits. The missing low-error points for LUCI were due to a lack of errors being found for those ensembles, while the missing point on the red curve in the rightmost plot is due to the threshold at that dropout rate being exceeded. Error bars are $5\sigma$ error bars using bootstrapping.
In the logical error per round heatmaps on the bottom row, LUCI outperforms the other two methods for every random dropout grid considered. For the same $p_{q}=p_{c}=0.01$ case, the improvement in median logical error rate per round is 25×25$\times$25 ×. he improvements in logical error per round follow from the distance improvements shown in the top set of heatmaps, and become even more significant at lower error rates. Consequently, the asymmetry between couplers and qubits becomes more appreciable as error rates reduce. An intuitive picture for this can be seen by considering the logical fault path, i.e. the logical operator found by adding the maxmimum likelihood matching from the correct and incorrect logical coset after decoding. This path is rarely exactly length $d$ for error rates near threshold, as the combinatoric coefficient in the logical error polynomial dominates and favors longer strings. However as physical error rates decrease the penalty for higher-weight strings begins to dominate and logical error rate performance is better described by distance, which is a worst-case metric for a quantum error correcting protocol.
@@
To further look at how LUCI helps when scaling quantum processors, we provide a set of teraquop footprint plots in Fig. 11. A teraquop footprint plot shows curves with logical error rate pinned at $10^{-12}$, hence the name. Along the x-axis we vary the physical error rate in the simulation, while the y-axis shows the physical footprint needed to attain the desired logical error rate. The three dropout methods are shown for $\left(p_{q},p_{c}\right)\in{(0.005,0.005),(0.01,0.01),(0.015,0.015),(0.02,0.02)}$, along with a baseline dropout-free circuit to act as a reference point. In the plot, we use error rates for a $\ell\times\ell\times\ell$ block for the regular surface code, a ××$\ell\times\ell\times 2\ell$ block for the Strikis/Brown and Auger/Bandage methods, and a ××$\ell\times\ell\times 4\ell$ block for the LUCI circuits, to account for the differences in timelike error. Despite this effectively doubling the logical error per round relative to the other methods, at $p_{q}=p_{c}=0.01$ and $p=0.001$ LUCI reduces the needed footprint by more than $25\%$ relative to the Auger/Bandage method. Further optimizations in constructing diagrams are possible to improve the timelike distance, and detailed stability experiments may allow us to shorted the timelike extent of the codeblocks.
IV Conclusion
@@
In this manuscript we presented LUCI, a framework for building error correction circuits adapted to dropout grids. The main benefits are its flexibility, and the reduction in penalty for broken qubits or couplers. For experimentally reasonable error rates and dropout parameters, we see improvements of 1.4×1.4$\times$1.4 × in distance and 25×25$\times$25 × in logical error rate per round. This improvement in logical error rate will become more significant as systems improve in both size and performance. The flexiblity of the method allows for swapping data and measure qubit roles, as well as dynamically editing QEC circuits to work around TLS’s that appear on the device. Our hope is that this reduction in the cost of dropouts enables superconducting qubit hardware research groups to further push the envelope during fabrication.
@@
While we do not discuss logical gates in this work, LUCI circuits should still admit the same operations as the standard surface code. In particular, the boundaries are the same as the usual surface code, so lattice surgery should be a viable option for logical Bell measurements and CNOT operations, but implementation is left to future work. There are also a number of possible performance optimizations, both for generating better LUCI diagrams and integrating known techniques in the field like shells, which are discussed in App. E. We have also seen that ensembled decoders such as Harmony Shutty et al. (2024) and Libra Jones (2024) perform well for LUCI circuits, indicating that there could be value in optimizing decoders for LUCI circuits.
@@
The flexibility provided by LUCI can also be used to build error correction circuits which are anisotropic and aperiodic, or even randomly generated. Along with the fact that LUCI circuits do not measure a consistent end-cycle error correcting code, this indicates that LUCI could be a possible step towards code-free fault-tolerant processes, like Nickerson and Bombín (2018); Newman et al. (2020); Hastings and Haah (2021); Fu and Gottesman (2024).
V Author Contributions
@@
VI Acknowledgements
The authors would like to thank Hartmut Neven, the Google Quantum AI team, and Google leadership, for the environment and resources to enable this research. The authors would also like to thank Dan Browne, Laleh Aghababaie Beni, Oscar Higgott, Matt Reagor, Dave Bacon, Austin Fowler, Alec Eickbusch, Alexis Morvan, Kevin Satzinger, and Cody Jones, for helpful input on the project and manuscript.
@@ Appendix A SI1000 Noise Model
The noise model used in this paper is SI1000, a superconducting inspired noise model which assumes a $1000ns$ cycle time. It is a single parameter noise model defined as follows:
@@
$\text{R}_{Z}(p)$ Initialize the qubit as |$\ket{0}$, followed by a bitflip channel of strength $p$.
$\text{R}_{X}(p)$ Initialize the qubit as $\ket{+}$, followed by a phaseflip channel of strength $p$.
@@
Table 1: Modified from Gidney et al. (2022). Noise channels and the rules used to apply them. Noisy rules stack with each other - for example, Idle($p$) and ResonatorIdle($p$) can both apply depolarizing channels in the same time step.
@@
Table 2: Modified from Gidney et al. (2022). Details of the error models used in this paper. See Table 1 for definitions of the noisy gates.
@@ Appendix B Dropouts near boundaries or clusters
Figure 12: An example situation in which a missing qubit near the boundary can lead to an unpairable gauge operator. The dotted red triangle indicates where a gauge would have to have existed to pair with the existing X-type gauge above the missing qubit. Since the dotted gauge is not present, there is no way to pair the X-type gauge into a super-stabilizer which commutes with its neighbors, so it must be removed.
Following the example in Ref Auger et al. (2017), dropouts near the boundary can lead to a gauge which does not have another gauge to join with to form a super-stabilizer, as seen in Fig. 12. In this case, there ends up being no solution to the pairing problem which includes this gauge operator, and it must be removed from the circuit. Once removed, the adjacent gauge operators can be individually promoted to stabilizers, as there no longer is an anticommuting operator present in the gauge group. This leads to distance only reducing in one basis. Note that in the example the remaining gauge operators are shown as triangles, which would be measured by L-shapes, but a possibly more performant solution would be to use I-shapes instead.
Another case where gauge operators end up tricky to pair is when there are multiple qubits broken near each other. In Auger et al. (2017) and Wei et al. (2024), they handle the case of diagonal broken data qubits by using so-called “bridge” qubits, where a measure qubit is used to measure a weight-2 gauge operator on diagonal data qubits. This weight-2 gauge commutes with the pairs of gauges on either side, allowing for two separate super-stabilizers of the opposite basis. In the mid-cycle picture LUCI operates in, measurements are done in place, so this disjoint weight-2 gauge operator would be split into two weight-1 measurements. To make the commutations work out, the two super-stabilizers of the opposite type which touch the weight-1 gauge operators must be merged, unlike in the end-cycle case.
Appendix C Data qubit dropout
Figure 13: The mid-cycle state for a grid with two qubit dropouts. The bottom left dropout creates Z(X)-type stabilizers which are extended vertically(horizontally). This means that the X(Z)-type logical operator, shown in red(blue), can use the extended stabilizer as a shortcut to reduce the logical distance by one. In contrast, the top right dropout has stabilizers extended parallel to the logical operator of their own type, so they do not reduce distance.
In the main text we mention that LUCI has spacelike distance reduced by 1 in both bases when a data qubit is broken, compared to 0 for a measure qubit. This is not due to the qubit roles, which in LUCI are arbitrary, but due to the orientations of the super-stabilizers formed to adapt the circuit to the missing qubit. In Fig. 13, we show the orientations of super-stabilizers in the mid-cycle state for a grid with two missing qubits. It can be seen that one qubit has super-stabilizers formed such that the X(Z)-type super-stabilizer is elongated in the direction of the Z(X)-type logical operator, reducing distance in both cases, while the other qubit has the super-stabilizers elongated perpendicular to the relevant logical operator. In the main text we explain this behavior using the language of data and measure qubits for simplicity, but the cause of this behavior is geometric, and independent on whether one compiles a circuit that would measure the missing qubit or not.
Appendix D Detector slices for LUCI example
@@
Appendix E Using Shells with LUCI
In the appendix of Wei et al. (2024) they discuss how their technique benefits from shells, a technique for dropouts first described in Ref. Strikis et al. (2023) which uses the schedule-induced gauge fixing technique introduced in Ref. Higgott and Breuckmann (2021). The basic idea is that by measuring gauge operators of the same type some number of times before switching bases, one can treat the gauge operators like stabilizers in the successive repetitions, with the super-stabilizer only being used in the first round of a given basis, as the individual gauge operators would be scrambled by the measurements of the opposite basis. In Ref. Strikis et al. (2023) they repeat the gauge operators around the dropout a number of times proportional to the patch diameter of the dropout region, while in Ref. Wei et al. (2024) they use global X and Z layers are study different ways to weight the patch diameters in the circuit.

### v2

@@ 3 Results
In this section we will compare LUCI to the two other methods of handling dropout mentioned in the introduction. We will refer to the method from [11] and [15] as the Auger/Bandage method, and the method from [12, 13, 14] as the Strikis/Brown method. As shown in [15], the Strikis/Brown method removes additional qubits and is strictly worse than the Auger/Bandage method, but is helpful because the single-type boundaries make certain proofs more straightforward. We keep it as a comparison point since it is well known, and as a result serves as a point of reference. We use stim for all simulations, with a correlated minimum-weight perfect matching decoder used for decoding and SI1000 noise [21], which sets error rates for different operations based on a single parameter $p$, described in Appendix A. We simulate X and Z logical memory experiments for $\ell\times\ell\times 4\ell$ code blocks, where $\ell$ is the patch diameter of the grid, and convert the resulting average logical error to a per-round logical error rate. For LUCI circuits we define a round as a single round of the LUCI diagram, meaning that we will use each round $\ell$ times over the course of the $4\ell$ rounds. As these different methods have different timelike distance, this comparison is slightly complicated, as depending on the timelike distance of your circuit you may need more rounds. LUCI circuits only promise to measure all stabilizers over four rounds, while the other two methods measure all stabilizers in two. We note, however, that the LUCI circuits still measure the majority of stabilizers at the same rate as standard methods, and only at low error rates will the sparse low-weight logical operators dominate performance. In practice, one should run stability experiments to chose the number of rounds in a logical idle block [22]. In the results presented in Figs. 10 and 11, we report per-round logical error rates, and a conservative reader may wish to multiply the logical error rate per round numbers by a factor of two to be consistent with the timelike distance of the other methods. The circuits used are provided on Zenodo[23].
@@
For the Auger/Bandage method and LUCI, we use custom written circuit generation tools, while for the Strikis/Brown method we use the circuit generation software helpfully provided in [14]. These methods are both amenable to using shells, first described in [24] as gauge-fixings, and then used in [12] to prove the existence of thresholds for surface codes under dropouts. In our comparison we do not use shells, but [15] discusses how they improve performance for the bandage method, and we expect the same would hold for LUCI. Appendix D discusses how one could implement shells with LUCI.
Figure 12: Teraquop footprint plot for LUCI. 100 grids were sampled for each error rate and dropout rate at $\ell\in{7,11,15}$ and then projected to $10^{-12}$. Error rates are for codeblocks of size $\ell\times\ell\times\ell$ for the dropout-free baseline circuits, $\ell\times\ell\times 2\ell$ for the Strikis/Brown and Auger/Bandage circuits, and $\ell\times\ell\times 4\ell$ for the LUCI circuits. The missing low-error points for LUCI were due to a lack of errors being found for those ensembles, while the missing point on the red curve in the rightmost plot is due to the threshold at that dropout rate being exceeded. Error bars are $5\sigma$ error bars using bootstrapping.
@@
In Fig. 10, we compare the three methods on the example grid presented in Fig. 1. The LUCI circuit used is the one built in Figs. 7 and 8. By looking at the slopes of the three curves, along with the gray reference curve, we can see the impact of distance loss as the error rate improves. The baseline distance-9 circuit has a slope of 5 in the log-log plot, because it takes $\lfloor\frac{d+1}{2}\rfloor$ errors to cause a logical error. When dropouts are added into the device, the distance is reduced to 3 for the Strikis/Brown and Auger/Bandage methods, and 7 for LUCI. The difference between the Strikis/Brown and Auger/Bandage methods is around the broken data qubit and broken coupler labeled as (e) in Fig. 1. As explained in Ref. [15], this configuration produces a “bridge” qubit, which the Auger/Bandage method preserves and the Strikis/Brown method removes. Otherwise the two circuits are very similar, and perform near identically as seen in the figure. LUCI also keeps this qubit, and only removes an additional qubit for the two-broken-coupler and broken-coupler-near-broken-qubit configurations labeled by (d) and (f) in Fig. 1. At an error rate of 0.001, LUCI shows more than a $30\times$ improvement in logical error rate per round over Auger/Bandage and Strikis/Brown methods for this example grid.
@@
In the logical error per round heatmaps on the bottom row, LUCI outperforms the other two methods for every random dropout grid considered. For the same $p_{q}=p_{c}=0.01$ case, the improvement in median logical error rate per round is $25\times$. The improvements in logical error per round follow from the distance improvements shown in the top set of heatmaps, and become even more significant at lower error rates. Consequently, the asymmetry between couplers and qubits becomes more appreciable as error rates reduce. An intuitive picture for this can be seen by considering the logical fault path, i.e. the logical operator found by adding the maximum likelihood matching from the correct and incorrect logical coset after decoding. This path is rarely exactly length $d$ for error rates near threshold, as the combinatoric coefficient in the logical error polynomial dominates and favors longer strings. However as physical error rates decrease, the penalty for higher-weight strings begins to dominate and logical error rate performance is better described by distance, which is a worst-case metric for a quantum error correcting protocol.
@@
To further look at how LUCI helps when scaling quantum processors, we provide a set of teraquop footprint plots in Fig. 12. A teraquop footprint plot shows curves with logical error rate pinned at $10^{-12}$, hence the name. Along the x-axis we vary the physical error rate in the simulation, while the y-axis shows the physical footprint needed to attain the desired logical error rate. The three dropout methods are shown for $\left(p_{q},p_{c}\right)\in{(0.005,0.005),(0.01,0.01),(0.015,0.015),(0.02,0.02)}$, along with a baseline dropout-free circuit to act as a reference point. In the plot, we use error rates for a $\ell\times\ell\times\ell$ block for the regular surface code, a $\ell\times\ell\times 2\ell$ block for the Strikis/Brown and Auger/Bandage methods, and a $\ell\times\ell\times 4\ell$ block for the LUCI circuits, to account for the differences in timelike error. Despite this effectively doubling the logical error per round relative to the other methods, at $p_{q}=p_{c}=0.01$ and $p=0.001$ LUCI reduces the needed footprint by more than $25\%$ relative to the Auger/Bandage method. Further optimizations in constructing diagrams are possible to improve the timelike distance, and detailed stability experiments may allow us to shorten the timelike extent of the codeblocks.
4 Conclusion
@@
In this manuscript we presented LUCI, a framework for building error correction circuits adapted to dropout grids. The main benefits are its flexibility, and the reduction in penalty for broken qubits or couplers. For experimentally reasonable error rates and dropout parameters, we see improvements of $1.4\times$ in distance and $25\times$ in logical error rate per round. This improvement in logical error rate will become more significant as systems improve in both size and performance. The flexibility of the method allows for swapping data and measure qubit roles, as well as dynamically editing QEC circuits to work around TLS’s that appear on the device. Our hope is that this reduction in the cost of dropouts enables superconducting qubit hardware research groups to further push the envelope during fabrication.
@@
While we do not discuss logical gates in this work, LUCI circuits should still admit the same operations as the standard surface code. In particular, the boundaries are the same as the usual surface code, so lattice surgery should be a viable option for logical Bell measurements and CNOT operations, but implementation is left to future work. There are also a number of possible performance optimizations, both for generating better LUCI diagrams and integrating known techniques in the field like shells, which are discussed in App. D. We have also seen that ensembled decoders such as Harmony [25] and Libra [26] perform well for LUCI circuits, indicating that there could be value in optimizing decoders for LUCI circuits.
@@
The flexibility provided by LUCI can also be used to build error correction circuits which are anisotropic and aperiodic, or even randomly generated. Along with the fact that LUCI circuits do not measure a consistent end-cycle error correcting code, this indicates that LUCI could be a possible step towards code-free fault-tolerant processes, like [27, 28, 29, 30].
5 Author Contributions
@@
6 Acknowledgements
The authors would like to thank Hartmut Neven, the Google Quantum AI team, and Google leadership, for the environment and resources to enable this research. The authors would also like to thank Dan Browne, Laleh Aghababaie Beni, Oscar Higgott, Matt Reagor, Dave Bacon, Austin Fowler, Alec Eickbusch, Alexis Morvan, Kevin Satzinger, Hao Tran and Cody Jones, for helpful input on the project and manuscript, and Joe Iverson and Catherine Leroux for pointing out an issue with one of the figures.
@@ Appendix A SI1000 Noise Model
The noise model used in this paper is SI1000, a superconducting inspired noise model which assumes a $1000ns$ cycle time. It is a single parameter noise model with noisy gates defined in Table 1 and error rates defined in Table 2:
@@
[…]
@@
Table 1: Modified from [31]. Noise channels and the rules used to apply them. Noisy rules stack with each other - for example, Idle($p$) and ResonatorIdle($p$) can both apply depolarizing channels in the same time step.
@@
Table 2: Modified from [31]. Details of the error models used in this paper. See Table 1 for definitions of the noisy gates.
@@ Appendix B Dropouts near boundaries or clusters
Figure 13: An example situation in which a missing qubit near the boundary can lead to an unpairable gauge operator. The dotted blue triangle indicates where a gauge would have to have existed to pair with the existing Z-type gauge above the missing qubit. Since the dotted gauge is not present, there is no way to pair the Z-type gauge into a super-stabilizer which commutes with its neighbors, so it must be removed.
Following the example in Ref [11], dropouts near the boundary can lead to a gauge which does not have another gauge to join with to form a super-stabilizer, as seen in Fig. 13. In this case, there ends up being no solution to the pairing problem which includes this gauge operator, and it must be removed from the circuit. Once removed, the adjacent gauge operators can be individually promoted to stabilizers, as there no longer is an anticommuting operator present in the gauge group. This leads to distance only reducing in one basis. Note that in the example the remaining gauge operators are shown as triangles, which would be measured by L-shapes, but a possibly more performant solution would be to use I-shapes instead.
Figure 14: Example of a clustered horizontal dropout configuration. Unlike a case where the middle two weight-1 X-type gauge operators are measured with a shared measure qubit, the four Z-type gauges are all combined into a single mid-cycle superstabilizer.
Another case where gauge operators end up tricky to pair is when there are multiple qubits broken near each other. In [11] and [15], they handle the case of diagonal broken data qubits by using so-called “bridge” qubits, where a measure qubit is used to measure a weight-2 gauge operator on diagonal data qubits. This weight-2 gauge commutes with the pairs of gauges on either side, allowing for two separate super-stabilizers of the opposite basis. In the mid-cycle picture LUCI operates in, measurements are done in place, so this disjoint weight-2 gauge operator would be split into two weight-1 measurements, as showing in Fig. 14. To make the commutations work out, the two super-stabilizers of the opposite type which touch the weight-1 gauge operators must be merged, unlike in the end-cycle case.
Appendix C Detector slices for LUCI example
@@
Appendix D Using Shells with LUCI
In the appendix of [15] they discuss how their technique benefits from shells, a technique for dropouts first described in Ref. [12] which uses the schedule-induced gauge fixing technique introduced in Ref. [24]. The basic idea is that by measuring gauge operators of the same type some number of times before switching bases, one can treat the gauge operators like stabilizers in the successive repetitions, with the super-stabilizer only being used in the first round of a given basis, as the individual gauge operators would be scrambled by the measurements of the opposite basis. In Ref. [12] they repeat the gauge operators around the dropout a number of times proportional to the patch diameter of the dropout region, while in Ref. [15] they use global X and Z layers are study different ways to weight the patch diameters in the circuit.